1. You want to make up a 400ml mixture made up of 25% of solution A and 75% of solution B.how many ml of A do you need?
2. Twenty percent (20%) of your population have brown eyes.if you have 3 450 people in your population, how many of them have brown eyes?
1
Expert's answer
2020-10-05T16:58:33-0400

(1) Amount of solution A needed"= \frac{25}{100} \times 400 ml = 100ml"


(2) Population with brown eyes "= \frac{20}{100} \times 3450 = 690" people